OpenStudy (anonymous):

write equation in slope-intercept form of line passing through (4,2) and (-3,1)

OpenStudy (anonymous):

find slope first: (2-1) / (4--3) = 1/7 then use y-y = m(x-x) point slope form, replace m with (1/7) and then the second y and x with either (4,2) or (-3,1). I choose (4,2): y-2 = (1/7)(x-4) distribute: y-2 = (1/7)x - 4/7 combine like terms to put into slope intercept format (y=mx+b): y = (1/7)x - 4/7 +2 y = (1/7)x - 4/7 +14/7 y = (1/7)x +10/7 :)